Journalist Report – November 13th

Journalist report;

Today we launched a rocket. Not just any rocket a repurposed Nike sounding rocket from the early days of the space program when we were still defining the winds in the upper atmosphere. At the launch we established effective range safety and cleared our airspace with an official NOTAM to the regulatory agencies controlling the airspace around our habitat. The rocket flew to 8200 feet and was safely recovered in a nearby field after a successful parachute recovery. No lives were lost and no animals were hurt in this demonstration so we count is a resounding success. To commemorate our success we ate a hearty meal of quiche and cinnamon rolls. After our launch we retired to the habitat for a customized exercise program, our traditional communal dinner and debriefings.

Sol Summary – November 13th

Crew 198 Sol Summary Report 13Nov2018

Sol: 3

Summary Title: A Successful Rocket Launch!

Author’s name: Jayne Paulson

Mission Status: Successful Launch and Retrieval of the MDRS Survey Rocket!

Sol Activity Summary: We learned about orbital mechanics and successfully launched a high powered rocket to 8,311 feet

Look Ahead Plan: It’s like Christmas everyday- we never know what it’s going to be!

Anomalies in work: None

Weather: Beautiful, sunny

Crew Physical Status: Crew passed Physical Readiness Test

EVA: Rocket Launch

Reports to be filed: Operations, Green Hab, Journalist, EVA Request, EVA Report, Sol Summary, Crew Bios

Support Requested: None at this time.

EVA Report – November 13th

EVA #2

We received our FAA clearance, called in our NOTAM, and successfully launched our rocket.

EVA Crew Members – All + MDRS Director (Approved by the Director to have all present to view the launch)

Location

Approximately 500m northeast of the Hab.

EVA Objectives

Rocket Launch

Results

We had a successful launch of the rocket. Initial estimates of an altitude are approximately 8,311 feet. We will have final data later today.

Issues/Mitigation

None

ATVs/Rovers Used

ATVs x 3, EVs x3

Operations Report – November 13th

Crew 198 Operations Report 13-Nov-2018

SOL: 3

Name of person filing report: Dana Levin

Non-nominal systems: Report filing, Rover Demos, engineering airlock electrical sockets, engineering airlock light, lower deck hab wall heater,

Notes on non-nominal systems:

– Generator update; after resetting the inverters to enable it to communicate load requirements to the generator generator seems to be responding to load needs and functioned without difficulty last night. It was started at 83% today and appears to be functioning well as of this report

– Report filing repeating from last night for broadcasting: we have had continued trouble with non updated report templates requiring a back and forth between MDRS and mission support demanding reports filed in different formats. This yielded significant crew frustration

o Proposed solution: reports could be turned into google forms. The links to these forms sent to the station crew to standardize filing format. This would also allow centralized data collection and presentation including some values that could be graphed over time and organized for presentation/analysis easily.

o We attempted to formulate such forms but our internet connection did not allow connection to the google servers

– Sojourner: Shannon had used Sojourner two days ago but the battery drained to the point where she was unable to safely return to the habitat. Rover was left about 4 miles north of the Hab beside brae highway.

o Our schedule was rearranged to enable 2 of our instructors to leave the hab to recover the rover in the Hab car while our students were undergoing lectures by a third instructor. This was completed without incident and luckily our pre-planned schedule was sufficient to absorb this unexpected non crew/non sim related contingency.

– Engineering air lock light and power sockets are still inoperable.

– Demos Rover remains with battery drained, water has been replaced and rover is hooked up to charge cable, will evaluate in morning

– 350 ATV is inoperable, likely with a dead battery that has yet to be replaced

Generator (hours run): 12

Solar— SOC % (Before generator is run at night): 83% (see above)

Diesel Reading – 60%

Propane Reading – not taken

Ethanol Free Gasoline – 13 gallons.

Water (auxillary tank) – 64 gallons

Water (static tank) – 350 gallons

Auxillary to Static tank transfer– yes

Gallons transferred: 30

Water in GreenHab – tank at 90%, unsure how many gallons this is

Water (loft) – Static to Loft Pump used – yes

Water Meter: unknown where this is located

Toilet tank emptied: yes

Deimos rover used: no

Hours:

Beginning charge: 0%

Ending charge:

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge: 100%

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 85

Currently charging: yes

Opportunity rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 87

Currently charging: yes

Curiosity rover used:

Hours: 1

Beginning charge: 100

Ending charge: 81

Currently charging: yes

Notes on rovers: (See above)

Demos: batteries refilled, attempting recharging overnight

Sojourner recovered and appears functional

ATV’s Used: (Honda, , 350.1, 350.2)

Reason for use: EVA

Oil Added? no

ATV Fuel Used: 1 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: one of the 350s is still non functional. The battery appears dead

HabCar used and why, where? No

CrewCar used and why, where? no

General notes and comments:

Please standardize reports and consider loading them into google forms or some kind of email template with a hab specific email. This saves data by preventing redownloading of mail to different accounts, avoids incorrect forms transmitted to crew or filed by crew, and makes data transfer easier

Summary of internet: functional
Summary of suits and radios: using our own
Summary of Hab operations: nominal despite facility difficulties Summary of GreenHab operations: none
Summary of ScienceDome operations: none
Summary of RAMM operations: none
Summary of any observatory issues: none identified as we are not using observatory
Summary of health and safety issues: none

Questions, concerns and requests to Mission Support: see above

Greenhab Report – November 13th

Crew 198 GreenHab Report – 13-NOV-2018

GreenHab Officer: Andrei

Environmental control: (Choose which of the following is appropriate and explain further if needed)

Heating

Shade cloth on

Average temperatures:

Low: 70

High 80

Hours of supplemental light: none

Daily water usage for crops: 3 gallons

Water in Blue Tank—(full) unknown number of gallons

Time(s) of watering for crops: 0900

Changes to crops: Note all emergence of seedlings, death of seedlings, etc (These are also mandatory to write in the GreenHab book for 2018-2019) none

Narrative: no other information at this time

Harvest: This is our research so it is important that all harvested crops are noted and also weighed and that weight recorded in the GreenHab book and here.

Support/supplies needed:none

operations report – november 12th

Crew 198 Operations Report 12-Nov-2018

SOL: 2

Name of person filing report: Dana Levin

Non-nominal systems: Generator, Report filing, Rover Sojourner, engineering airlock electrical sockets, engineering airlock light, lower deck hab wall heater,

Notes on non-nominal systems:

– Generator: Last night the generator had an underspeed cutout failure. While our crew has not been trained on generator mantinance and repair we were able to obtain assistance from our local benevolent Martian overlord Shannon.

– Event log;

o Around 1930 generator started with SOC at 46%

o 2000 generator cutout, temperature reading at 212F reporting a generator underspeed alert, generator restarted

o 2100 generator failed again, restarted Shannon alerted

o 2140 generator failed again and inverters cutout killing power to the entire MDRS complex. We drained coolant from the overflow tank including removing some congealed substance off the top of the tank, generator left off overnight

o At 0645 SOC was 26% solar began charging batteries

o 0800 generator started after evaluation by DG at 1800 RPM, continued operation for 1 hour and was shut off

o TODAY: 1720 SOC at 84% generator started as sunset. 1745 generator cutoff again.

– Current theory; mechanical obstruction to generator prevents full RPM which means current ops normal habitat power load exceeds generation capacity and generator stalls out

– Report filing: we have had continued trouble with non updated report templates requiring a back and forth between MDRS and mission support demanding reports filed in different formats. This yielded significant crew frustration

o Proposed solution: reports could be turned into google forms. The links to these forms sent to the station crew to standardize filing format. This would also allow centralized data collection and presentation including some values that could be graphed over time and organized for presentation/analysis easily.

o We attempted to formulate such forms but our internet connection did not allow connection to the google servers

– Sojourner: Shannon had used Sojourner two days ago but the battery drained to the point where she was unable to safely return to the habitat. Rover was left about 4 miles north of the Hab beside brae highway.

o Our schedule was rearranged to enable 2 of our instructors to leave the hab to recover the rover in the Hab car while our students were undergoing lectures by a third instructor. This was completed without incident and luckily our pre-planned schedule was sufficient to absorb this unexpected non crew/non sim related contingency.

– Engineering air lock light and power sockets are still inoperable.

– Demos Rover remains with battery drained and empty of water, rover is inoperable.

– 350 ATV is inoperable, likely with a dead battery that has yet to be replaced

Generator (hours run): 2

Solar— SOC % (Before generator is run at night): 84% (see above)

Diesel Reading – 60%

Propane Reading – not taken

Ethanol Free Gasoline – 14 gallons.

Water (auxillary tank) – 60 gallons

Water (static tank) – 380 gallons

Auxillary to Static tank transfer– yes

Gallons transferred: 20

Water in GreenHab – full tank, unsure how many gallons this is

Water (loft) – Static to Loft Pump used – yes

Water Meter: unknown where this is located

Toilet tank emptied: no

Deimos rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: recovered and currently charging

Hours: 1 hour

Beginning charge: 31%

Ending charge: 40% (towed)

Currently charging: yes

Spirit rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 89

Currently charging: yes

Opportunity rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 87

Currently charging: yes

Curiosity rover used:

Hours: 1

Beginning charge: 100

Ending charge: 81

Currently charging: yes

Notes on rovers: (See above)

ATV’s Used: (Honda, , 350.1, 350.2)

Reason for use: EVA

Oil Added? no

ATV Fuel Used: 1 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: one of the 350s is still non functional. The battery appears dead

HabCar used and why, where? Yes, to retrieve sojourner from Brae highway and to pump water into greenhab

CrewCar used and why, where? no

General notes and comments:

Please standardize reports and consider loading them into google forms or some kind of email template with a hab specific email. This saves data by preventing redownloading of mail to different accounts, avoids incorrect forms transmitted to crew or filed by crew, and makes data transfer easier

Summary of internet: functional

Summary of suits and radios: using our own

Summary of Hab operations: nominal despite facility difficulties

Summary of GreenHab operations: none

Summary of ScienceDome operations: none

Summary of RAMM operations: none

Summary of any observatory issues: none identified as we are not using observatory

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support: see above

journalist report – november 12th

Today was our first full day on Mars. We started with very limited power thanks to the system failures detailed by Shannon previously. We also had some lectures from Rick on various space medicine topics, including ultrasound in space flight, winter survival, water survival, and the role of the flight surgeon. We also conducted an EVA and simulated a medical contingency while descending rocky scree. Finally, we had a lecture from Sean Serell, our rocketry specialist, who is directing the launching of our rocket tomorrow. The crew is tired in a good way, and very excited for tomorrow!

EVA report – november 12th

Crew 198 EVA Report 12-Nov-2018

EVA #2

Author: Dana Levin

Purpose of EVA: medical training, establishing com relay

Start time: 1300

End time: 1530

Narrative: relay established medical contingencies handles

Destination: Kissing Camel ridge west

Coordinates (use UTM NAD27 CONUS):

Easting: 518300

Northing: 4249300
Participants: Dan, Joe, Chuck, Jayne, Bindhu, Aakar, Andrei, Ben, Andy

Road(s) and routes per MDRS Map Cowdung Road

Mode of travel:

ATVs x 3 (specifically approved by MDRS Director), EVs x3

greenhab report – november 12th

Crew 198 GreenHab Report – 12-NOV-2018

GreenHab Officer: Andrei

Environmental control: (Choose which of the following is appropriate and explain further if needed)

Heating

Shade cloth on

Average temperatures:

Low: 70

High 80

Hours of supplemental light: none

Daily water usage for crops: 4 gallons

Water in Blue Tank—(full) unknown number of gallons

Time(s) of watering for crops: 0900

Changes to crops: Note all emergence of seedlings, death of seedlings, etc (These are also mandatory to write in the GreenHab book for 2018-2019) none

Narrative: no other information at this time

Harvest: This is our research so it is important that all harvested crops are noted and also weighed and that weight recorded in the GreenHab book and here.

Support/supplies needed:none